Can I please get the steps using a BA11 Plus calculator for problems 21 and 22?

c. $7,324.89 d. $7,011.87 e. $7,854.13 21. You are buying your first house for $220,000, and are paying $30,000 as a down payment. You have arranged to finance the remaining $190,000 30-year mortgage with a 7% nominal interest rate and monthly payments. What are the equal monthly payments you must make? a. $1,513 b. $1,110 c. $1,264 구), 12-0593-江 140,000 p,V e. $1,349 22. Suppose you borrowed $25,000 at a rte of8% and must repay it in 4 equal installments at the end of each of the next 4 years. How much would you still owe at the end of the first year, after you have made the first payment? a. $21,110 b. $19,452 c. $18,513 d. $18,976 e. $19,049 You just deposited $5,000 in a bank account that pays a 12% nonínal interest rate. conipotided 23. you also add another $10,000 to the account one year (12 months) from now and another S15,000 to the account two years from now. how much will be in the account three years (36 months) from now? a. $36,753.57 b. $33.431.83 c. $34.764.40 d. $37,843.15 e. $35,I19.76 NA 24. Which of the following is NOT CORRECT?
Answer #1

21) Calculation of monthly Payment: Monthly payment can be calculated using PMT(RATE,NPER,PV,FV,TYPE) function in Excel as follows: Price of house Down-payment Loan taken Duration of mortgage Interest rate $220,000.00 30000 190,000 30 Years 7.00% 12 13 14 15 16 17 18 Monthly Payment can be calculated as below: RATE (Montly interst rate): NPER (No of Months): PV (Loan Amount): 0.58%-D10/12 360 D9*12 190,000-D8 (Negative as it is the amount received) TYPE Monthly Payment $1,264.07 PMT(D13,D14,D15,D16,D17) Hence Monthly Payment is Thus option (c) is correct. $1,264.07 In BA 11 Plus Calculator following data can be used to find the answer: 24 25 26 27 28 29 30 360 0.583% $190,000 PV $25,000.00 Loan Amount Duration of loan Interest rate Using following input in BA 11 plus calculator, PMT can be calculated: I (Annual interst rate): N(No of Payments) PV (Loan Amount): 32 4 Years 8% 34 35 36 37 38 39 8.00% $25,000 41 Annual Payment $7,548.02 PMT(D36,D37,D38,D39) 42 Mortgage table can be prepared as follows: Year Beginning BalandAnnual PaymentInterest Paid Principle Paid Ending Balance $25,000 $19,452 $13,460 $6,989 $7,548.02 $7,548.02 $7,548.02 $7,548.02 2,000.00 1,556.16 $1,076.81 $559.11 $5,548.02 $5,991.86 $6,471.21 $6,988.91 $19,451.98 $13,460.12 $6,988.91 $0.00 45 47 49 Thus amount owed at the end of first year is Thus the option (b) is correct. $19,452 52Formula sheet

A B C D E F G H I
2
3 21)
4 Calculation of monthly Payment:
5 Monthly payment can be calculated using PMT(RATE,NPER,PV,FV,TYPE) function in Excel as follows:
6 Price of house 220000
7 Down-payment 30000
8 Loan taken =D6-D7
9 Duration of mortgage 30 Years
10 Interest rate 0.07
11
12 Monthly Payment can be calculated as below:
13 RATE (Montly interst rate): =D10/12 =D10/12
14 NPER (No of Months): =D9*12 =D9*12
15 PV (Loan Amount): =-D8 =-D8 (Negative as it is the amount received)
16 FV 0
17 TYPE 0
18 Monthly Payment =PMT(D13,D14,D15,D16,D17) =PMT(D13,D14,D15,D16,D17)
19
20 Hence Monthly Payment is =D18
21 Thus option (c) is correct.
22
23 In BA 11 Plus Calculator following data can be used to find the answer:
24 N =D14
25 I =D13
26 PV =D15
27 FV =D16
28
29
30 22)
31
32 Loan Amount 25000
33 Duration of loan 4 Years
34 Interest rate 0.08
35 Using following input in BA 11 plus calculator, PMT can be calculated:
36 I (Annual interst rate): =D34
37 N(No of Payments): =D33
38 PV (Loan Amount): =-D32
39 FV 0
40
41 Annual Payment : =PMT(D36,D37,D38,D39) =PMT(D36,D37,D38,D39)
42
43 Mortgage table can be prepared as follows:
44 Year Beginning Balance Annual Payment Interest Paid Principle Paid Ending Balance
45 1 =D32 =$D$41 =D45*$D$36 =E45-F45 =D45-G45
46 =C45+1 =H45 =$D$41 =D46*$D$36 =E46-F46 =D46-G46
47 =C46+1 =H46 =$D$41 =D47*$D$36 =E47-F47 =D47-G47
48 =C47+1 =H47 =$D$41 =D48*$D$36 =E48-F48 =D48-G48
49
50 Thus the amount owed at the end of first year is =H45
51 Thus the option (b) is correct.
52